Twenty new members are elected to an agency every two years. The agency has 80 members on average overall. How long does a member hold his/her position at the agency?

A. 4 years

B. 10 years

C. 8 years

D. 5 years
Answer #1

Average yearly throughput of members = 20 / 2 years = 10 / year

As per Little Law, Time in a system , W = Average members in the system / Average throughput of members = 80/10 = 8 years